(a) If year 2 is the base year, give the price index for year 3.

                 (b) Give the nominal GDP for year 4. $120

                 (c) What is the real GDP for year 4?

                 (d) Tell which years you would deflate nominal GDP and which years you would inflate nominal GDP in finding real GDP.


1
Expert's answer
2021-05-25T16:56:56-0400

Solution:

a.). Price index for year 3 = Price  of  the  market  basket  in  current  yearPrice  of  the  market  basket  in  base  year×100\frac{Price\; of\; the\; market\; basket\; in \; current \; year}{Price\; of\; the\; market\; basket\; in \; base \; year} \times 100

Assume:

Price of the market basket in year 3 = 120

Price of the market basket in year 2 = 90

Price index for year 3 = 12090×100=133.33%\frac{120}{90}\times 100 = 133.33\%

 

b.). Nominal GDP = Current year quantity output x current year market price

Assume:

Current year output = 150

Nominal GDP = 150 ×\times 120 = 18,000

 

c.). Real GDP = Current year quantity output x base year market price

Assume:

Base year market price = 100

Real GDP for year 4 = 150 ×\times 100 = 15,000

 

d.). The year that would deflate nominal GDP is year 2, while the year that would inflate nominal GDP in finding real GDP is year 3.